All satellite details were emailed out this morning; please check!
Follow CBS_Satfacts on Twitter
My Tweets-
Recent Posts
Previous Productions
Search Content
Meta
Previous Production’s Info
All satellite details were emailed out this morning; please check!
We lost the signal for a few seconds at 2102GMT/2202CET on the satellite from New York.
We are now on the reserve feed.
Apologies to all
A reminder that Testing starts at 1545GMT/1645CET on IS-1002 (with an 18MHz English feed, and a German feed), on Telstar12 (English) and on AB3 with Spanish and French. (The 27MHz more powerful English feed on IS-1002 starts at 1745CET)
If you think there is a non-local problem with your reception, call the Helpline.
Hope you all enjoy this afternoon and evening.
Chris.
Services established on schedule.
Vision disturbances at 15:58GMT whilst we reset screen subtitle insertion units – all fine now.
27Mhz service now established and stable. (16:50GMT).
Service loss at 21:02GMT – switched to reserve connectivity – all stable again
Services ceased at 21:48 – We appologise for the break in transmission tonight
Good morning and a Happy New Year to you all
A reminder that there is a test on IS10-02 and AB3 at 0930GMT/1030CET this morning.
If you have any queries, call the Helpline.
I can confirm that the Test is now underway, until 1100GMT/1200CET.
The audio is VERY low level, is only of one microphone in an empty hall (you can hear fan-rumble), and there is nothing in the surround.
The video colour bars are coming from the uplink truck (not the OB van) and bear no relation to the sync-clap you can see at the bottom of the frame.
I am sure that this test is of very little use other than to tune in, and so please do come back at 1630GMT/1730CET when I hope we can get some professional service from the Event.
We are monitoring this feed for Arqiva, and providing reception advice.
I understand that the Test at 1000GMT/1100CET will be of colour bars and tone only. This is because the OB van in Berlin is unmanned at this time, and it was left this way last night.
Come back here to follow any developments
(All content today is outside of our control or influence)
I think that almost everyone of you had some issues when the broadcast went between the Test loop and the DCP-type file of the “Nutcracker”. We will be asking the broadcaster and other key-stakeholders to give us a report as to what happened today.
Please do come back here, where we will report any significant findings that are made known to us.
We are as disappointed as you.
Testing begins shortly for this afternoon’s Bolshoi transmission. Signals due up at 13:00GMT / 14:00CET.
At 14:46GMT/15:46CET, there will be a slight “bump” in the audio and video as we switch between the Test source and the “Nutcracker” source.
AB3 established and stable at 12:54GMT
10-02 established and stable at 12:56GMT
Vision bump at 14:45 resulted in loss of service – tuned away and then back to Bolshoi channel to regain vision and audio.
If your receiver has crashed and will not lock on following a channel change. Please do a full retune.
Services ceased at 17:15GMT.
We apologise for todays difficulties which were network wide. We are investigating the causes and particularly why the Icecrypt exhibited the effects it did. Any further reports on your experiences would be helpful (via email).
This is now well underway, and the signals are really HOT!
Hope it is going well for you
There is a Test for Sunday’s “Nutcracker” at 1000GMT/1100CET today.
There is a feed on IS-1002 and AB3
Please make the best use of this opportunity.